OpenStudy (anonymous):

Graph 3/x^2-4. Then find the domain/range/VA

OpenStudy (astrophysics):

\[f(x) = \frac{ 3 }{ x^2 }-4?\] is this your equation, think of the parent function, which in this case is \[\frac{ 1 }{ x^2 }\] it looks like  you can get this by simply making a table of values which you can do yourself. Once you have your parent function now we can do our transformations. \[y = \frac{ 3 }{ x^2 }\] where the 3 should give you a scaling factor then afterwards add the -4, \[y = \frac{ 3 }{ x^2 }-4\] where the -4 should give you a vertical translation of -4 units. If you're not comfortable at seeing it directly you can always make a table of values for each step, good luck!